Develop an activity that explains the concept of pathogen theorem.
1
Expert's answer
2020-06-04T21:22:32-0400

1 STEP : We need to make sure the formula



"\left(a+b\right)^2=a^2+2ab+b^2"

We will use the area method. Without proof, we believe that we understand the formulas for the area of a rectangle with sides "a" and "b" - "S=ab" , and the area of a square with side "a" - "S=a^2"

Then,


"S=\left(a+b\right)^2-\text{large square formula}"

This square consists of two rectangles and two different squares. Then,



"\boxed{\left(a+b\right)^2=a^2+2\cdot ab+b^2}"

Q.E.D.

2 STEP : We pass to the proof of the Pythagorean theorem.

We will also use the area method. We construct a large square with side (a + b) of four rectangular triangles as shown in the figure






"S=\left(a+b\right)^2=a^2+2ab+b^2-\text{large square formula}"

This square consists of four rectangular triangles with legs "a" and "b" and a square with side "c" , therefore



"a^2+2ab+b^2=c^2+4\cdot\frac{1}{2}\cdot a\cdot b\\[0.3cm]
a^2+2ab+b^2=c^2+2\cdot a\cdot b\\[0.3cm]
\boxed{a^2+b^2=c^2}"

Q.E.D.